5. Individual Problems

4-5 Your insurance firm processes claims through its two facilities: facility A and facility B. Each month, facility A handles 9,000 claims and incurs in $162,000 fixed costs and $180,000 in variable costs. Each month, facility B handles 11,000 and incurs $95,000 in fixed costs and $143,000 in variable costs.

Hint: Be careful not to commit the sunk-cost fallacy in your analysis.

If you anticipate a moderate decrease in the number of claims, you should lay off workers in facility A or B in order to decrease costs.
Answer #1

Average variable cost in A=180000/9000=20

Average variable cost in B=143000/11000=13

Since average variable cost is lower in Facility B then workers should be laid off from facility A.

Firm should Layoff from A in order to reduce/decrease cost